OpenStudy (anonymous):

Solve the system 3x - 5y = 8 and 4x - 7y = 10. A. (1, -1) B. (2, 6) C. (6, 2) D. no solution E. infinitely many solutions

OpenStudy (anonymous):

c

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

(3x - 5y = 8) * -4 (4x - 7y = 10) *3 -12x + 20y = -32 12x - 21y = 30 -1y=-2 y=2 then when y=2 3x - 10 = 8 x = 6 then the solution is C (6, 2)

OpenStudy (anonymous):

Why did you multiply is by -4?

OpenStudy (anonymous):

to eliminate the "x"

OpenStudy (anonymous):

But why by -4..?

OpenStudy (anonymous):

@Feel-Like-A-Sir to make the both equations has the same value of "x" .. one is positive and the other is negative .. so we can eliminate "x" and keep "y"

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

Alright, I understand it now. Thanks!